# PhyWorldBench: A Comprehensive Evaluation of Physical Realism in Text-to-Video Models

### v1

@@ 4.6 Breakdown in Physics at Higher Complexity
While many models handle simple, isolated physics well, their performance deteriorates under complex interactions involving multiple forces or constraints. For instance, in billiard collisions, models often fail to preserve momentum and friction, leading to unnatural stopping, incorrect deflections, or interpenetrating objects. Similarly, fluid simulations frequently lack realistic turbulence and surface tension when water interacts with solid surfaces. As demonstrated in Appendix M, prompts involving multiple objects introduce significantly higher requirements for semantic adherence. For instance, when tested with the prompt “A feather and a stone are dropped in the air, where both fall, and the stone drops much faster”, all models fail to represent the phenomenon correctly. Beyond failing to adhere to semantics, these errors highlight a broader issue: models struggle to depict the correct physical relationships between interacting objects, often treating them as independent entities rather than systems governed by unified physical laws. This suggests that while models have learned pattern-based approximations of physics, they lack a robust understanding of how multiple physical forces interact dynamically in complex environments.
@@
A major challenge in video generation is the tendency for models to prioritize cinematic aesthetics over strict physical realism. Well-performed models like Sora, Gen-3, and Pika often introduce exaggerated motion dynamics, such as objects moving with heightened fluidity or unnatural acceleration, making scenes appear choreographed rather than physically grounded. For example, an object that should fall naturally under gravity might instead descend too smoothly or float unrealistically, breaking expected motion physics. This issue is particularly evident in action sequences, where momentum conservation and force dynamics are distorted for dramatic effect, leading to visually striking but physically implausible results. The first example in Figure 10 presented an apple floating in the air, then it suddenly dropped as the camera moved; such stylized rendering, while beneficial for artistic storytelling, poses significant challenges for applications requiring strict physical fidelity.
@@ 5 Contributions and Future Work
We propose PhyWorldBench, a large-scale, thorough, and multi-dimensional benchmark for evaluating text-to-video generation models across diverse physics categories. It provides insights into which physical phenomena are hardest to simulate and what capabilities current models lack. Additionally, we propose an automated evaluator to measure physical accuracy and a straightforward receipt for designing prompt. Future work will involve continuously evaluating emerging video generation models, updating our benchmark to reflect advancements in the field, and refining our leaderboard and analysis to track progress in physical accuracy and model capabilities.
Impact Statement
Appendix A Benchmark Details
@@
Following  (Bansal et al., 2024) and  (Meng et al., 2024), we use Semantic Adherence (SA) and Physical Commonsense (PC) as metric to evaluate the video performance. SA measures if a caption aligns with video frames, while PC assesses if actions obey real-world physics. Annotated as $\text{SA},\text{PC}\in\{0,1\}SA , PC ∈ { 0 , 1 }$, where 1 indicates proper grounding. We report the fraction of videos satisfying $\text{SA}=1SA = 1$, $\text{PC}=1PC = 1$, and both jointly. We create SA and PC based on a thorough review of all prompts and extensive feedback from human annotators, we establish a set of evaluation criteria to systematically assess the generated videos, as illustrated in Figure 4. The evaluation is structured around two key standards, each encompassing specific rules:
@@
Appendix B Annotation Details
Prompt Annotation After designing the physics categories, we further prepared prompts within each category. To achieve this, we gathered undergraduate and graduate students majoring in Computer Science for an annotation session. They were instructed to provide high-quality annotations by creating both an Event Prompt and a corresponding Physics-enhanced Prompt that aligned with the primary physics category and sub-category. The Detailed Prompt was generated using GPT-4o with the MLLM prompt (as described in Table 11) and subsequently verified by human annotators. Finally, each prompt underwent a final verification by the authors.
@@ Appendix B Annotation Details
Appendix C Detailed Analysis on Evaluator
While some existing physics metrics have been proposed, none fit PhyWorldBench’s needs due to domain or structural mismatches (Guo et al., 2025; Zhang et al., 2025). For example, PhyGenEval (Meng et al., 2024) assumes a fixed event order, which many of PhyWorldBench’s scenarios do not have, and VideoCon‑Physics (Bansal et al., 2024) is trained specifically on its own VideoPhy benchmark, yielding sub‑optimal performance on PhyWorldBench—so we exclude both to avoid unfair comparisons.
In Table 6, we therefore evaluate our CAP across multiple MLLMs, including proprietary GPT‑4o, GPT‑o1, Gemini‑Flash‑2.0, and open‑source Qwen‑VL‑2.0. Crucially, CAP uses a two‑step chain‑of‑thought prompt—first asking the model to describe the video, then to reason about any physics issues—so it does not assume videos are incorrect by default and yields only minimal false‑positive increases (0.001 in Semantic Adherence, 0.014 in Physical Commonsense), confirming its neutrality. Moreover, because it operates solely on the final video and its physics‑based standards, CAP is agnostic to the underlying generative pipeline—whether text‑to‑video, image‑to‑video, video‑to‑video, or hybrid.
@@ Appendix C Detailed Analysis on Evaluator
Appendix D Model Application Details
@@
Appendix E Insights into Performance Between Models
Our analysis reveals that both model size and sampling budget significantly impact physics performance. Larger models consistently outperform smaller ones on foundational physics tasks, as shown in Table 7. For instance, Cosmos 14B outperforms Cosmos 7B across all SA/PC metrics, and CogVideo‑5B beats CogVideo‑1.5B, confirming that scaling yields clear gains in semantic adherence and physical commonsense.
@@ Appendix E Insights into Performance Between Models
Cosmos 14B (Agarwal et al., 2025) 0.374 0.268 0.212 0.331 0.243 0.182 0.091 0.069 0.020 0.333 0.241 0.184
Cosmos 7B (Agarwal et al., 2025) 0.354 0.243 0.201 0.339 0.221 0.185 0.088 0.054 0.020 0.323 0.218 0.178
CogVideo‑5B (Hong et al., 2022) 0.419 0.266 0.192 0.308 0.207 0.154 0.062 0.021 0.000 0.351 0.225 0.163
CogVideo‑1.5B (Hong et al., 2022) 0.382 0.247 0.175 0.246 0.199 0.148 0.073 0.021 0.010 0.310 0.210 0.150
@@
Appendix F Detailed Performance over All Physics Categories
In the main paper, we presented results across different levels of physics. Here, we provide a detailed breakdown of results for each specific physics category in Table 8.
@@ Appendix F Detailed Performance over All Physics Categories
CogVideo 5b 0.476 0.272 0.223 0.333 0.125 0.115 0.438 0.198 0.156 0.470 0.500 0.350 0.495 0.326 0.179 0.304 0.174 0.130 0.143 0.010 0.000 0.367 0.306 0.235 0.413 0.304 0.228 0.062 0.021 0.000
Step-video-T2V 0.401 0.265 0.204 0.273 0.202 0.166 0.361 0.239 0.145 0.371 0.264 0.227 0.384 0.273 0.238 0.328 0.227 0.154 0.165 0.000 0.000 0.388 0.385 0.25107 0.446 0.368 0.31466 0.085 0.021 0.020
@@ Appendix G Limitations
Appendix H Code and Data Availability
@@
  • Website: Additionally, this work is available on the NVIDIA Deep Imagination Research website at https://research.nvidia.com/labs/dir/phyworldbench/.
Appendix I Query Used
Here we list the three MLLM queries used in this work, corresponding to the first-stage description prompt (Table 9), the second-stage structured-answer prompt (Table 10), and the detailed-prompt refinement query (Table 11). Note that for proprietary MLLMs such as GPT-o1, GPT-4o and Gemini-2.0-Flash, the model may refuse to evaluate a small fraction of generated videos; the refusal rate is below 3
@@ Appendix I Query Used
Appendix J Qualitative Analysis on Individual Model
@@
  • CogVideo: Tends to produce distorted objects, such as non-spherical balls that appear dented. Movements are often excessively fast, blurred, erratic, or jittery, sometimes appearing unnaturally sped up.
@@
Figure 15: Physics Violations Observed in Videos Generated by the CogVideo 5b Model. a. Raindrops do not return to the clouds after hitting the ground; instead, they exhibit normal rainfall, failing to demonstrate anti-physics. b. Toy car and real car rolling down ramps are depicted as the same size, violating proportional realism. c. Car driving on an icy road shows no slippage, unrealistically maintaining normal traction. d. Basketball dropped onto the ground displays unnatural bouncing motion. e. Person standing on one leg during yoga exhibits unrealistic human anatomy.
@@
Appendix K Further Analysis on Generative Models Struggling with Certain Physical Scenarios
@@
Appendix L Leaderboard
Table 12 presents the leaderboard performance on both human evaluation and our designed CAP. The performance shows a highly consistency and only difference is that for proprietary models, Sora ranked number 2 while Kling ranked number 2 in CAP. Through further analysis, we found that Kling’s more cinematic style can bias CAP toward higher scores despite only moderate physical correctness, whereas our human annotators—who were instructed to focus strictly on physics plausibility—penalize these artifacts, causing Sora to rank higher under human evaluation.
@@
Appendix M Influence of Multiple Objects
@@
Appendix N Dataset Structure
@@
Appendix O Data Release
We publicly released a comprehensive dataset that includes the core assets of PhyWorldBench, the 1,050 JSON prompt files, the evaluation standards, and the physics categories and subcategories, used to assess physical realism in text-to-video models. We have also attached over 10k of the generated videos that we used for experimentation. The licensing terms for the videos will follow those set by the respective model creators, as referenced in this work, while the curated captions and human annotations will be provided under the MIT License. Additionally, our release will include standardized evaluation protocols, and evaluation scripts to facilitate rigorous assessment. The entire project will be open-sourced, ensuring free access for research and academic purposes.

### v2

@@ 4.6 Breakdown in Physics at Higher Complexity
Models’ performance deteriorates under complex interactions involving multiple forces or constraints. As demonstrated in Appx. R, prompts involving multiple objects introduce significantly higher requirements for semantic adherence. For instance, when with the prompt “A feather and a stone are dropped in the air, where both fall, and the stone drops much faster”, all models fail to present the phenomenon correctly. The current models lack a robust understanding of how multiple physical forces interact in complex environments. We similarly built 20 prompt pairs focused on interactions of multiple forces or collisions, applied them to the 12 models, and highlighting Pika as the best model as in Tab. 17. Physics further breaks down at higher complexity — 12% success overall vs. 18% for simpler cases; Pika 15% vs. 40%. This suggest that the current video models do not really understand the physics thoroughly and are more likely to mimic the pattern in the training set.
@@
A major challenge in video generation is the tendency for models to prioritize cinematic aesthetics over strict physical realism. Well-performed models like Sora, Gen-3, and Pika often introduce exaggerated motion dynamics, such as objects moving with heightened fluidity or unnatural acceleration, making scenes appear choreographed rather than physically grounded. For example, an object that should fall naturally under gravity might instead descend too smoothly or float unrealistically. This issue is particularly evident when momentum conservation and force dynamics are distorted for dramatic effect. The first example in Figure 11 presented an apple floating in the air, then it suddenly dropped as the camera moved; such stylized rendering, while beneficial for artistic storytelling, poses significant challenges for applications requiring strict physical fidelity.
@@ 5 Contributions and Future Work
We propose PhyWorldBench, a large-scale, thorough, and multi-dimensional benchmark for evaluating text-to-video generation models. It provides insights into which physical phenomena are hardest to simulate and what capabilities current models lack. Additionally, we propose an automated evaluator to measure physical accuracy and a straightforward receipt for designing prompt. Future work will involve continuously evaluating emerging video generation models, updating our benchmark to reflect advancements in the field, and refining our leaderboard and analysis to track progress in physical accuracy and model capabilities. LLM usage is described in Appx. A.
Ethics statement
The authors confirm that we have carefully reviewed and adhered to the ICLR Code of Ethics in this work.
Reproducibility statement
We will open-source our codebase to ensure reproducibility. We will first release all prompts, evaluation standards, and We will release the full evaluation code of our benchmark. We will also release all the generated videos and their corresponding prompts.
Appendix A LLM Usage Statement
The large language model is used during the manual script proofreading to detect grammatical errors and rephrase some overly long sentences. Meanwhile, as mentioned in the paper, our designed evaluator is based on LLM, the prompt draft creation process also involved LLM.
Appendix B Benchmark Details
@@
Following  Bansal et al. (2024) and  Meng et al. (2024), we use Semantic Adherence (SA) and Physical Commonsense (PC) as metric to evaluate the video performance. SA measures if a caption aligns with video frames, while PC assesses if actions obey real-world physics. Annotated as $\text{SA},\text{PC}\in\{0,1\}$, where 1 indicates proper grounding. We report the fraction of videos satisfying $\text{SA}=1$, $\text{PC}=1$, and both jointly. We create SA and PC based on a thorough review of all prompts and extensive feedback from human annotators, we establish a set of evaluation criteria to systematically assess the generated videos. The evaluation is structured around two key standards, each encompassing specific rules:
@@
Appendix C Annotation Details
Prompt Annotation After designing the physics categories, we further prepared prompts within each category. To achieve this, we gathered undergraduate and graduate students majoring in Computer Science for an annotation session. They were instructed to provide high-quality annotations by creating both an Event Prompt and a corresponding Physics-enhanced Prompt that aligned with the primary physics category and sub-category. The Detailed Prompt was generated using GPT-4o with the MLLM prompt (as described in Tab. 12) and subsequently verified by human annotators. Finally, each prompt underwent a final verification by the authors.
@@ Appendix C Annotation Details
Appendix D Detailed Analysis on Evaluator
While some existing physics metrics have been proposed, none fit PhyWorldBench’s needs due to domain or structural mismatches (Guo et al., 2025; Zhang et al., 2025a). For example, PhyGenEval (Meng et al., 2024) assumes a fixed event order, which many of PhyWorldBench’s scenarios do not have, and VideoCon‑Physics (Bansal et al., 2024) is trained specifically on its own VideoPhy benchmark, yielding sub‑optimal performance on PhyWorldBench—so we exclude both to avoid unfair comparisons. Besides, VideoPhy-2 (Bansal et al., 2025) focused on action and dynamic instead of general physics, and its evaluator is also designed and finetuned on the associated benchmark annotation, therefore for the same reason, we also did not include it in this paper.
In Tab. 6, we therefore evaluate our CAP across multiple MLLMs, including proprietary GPT‑4o, GPT‑o1, Gemini‑Flash‑2.0, and open‑source Qwen‑VL‑2.0. Crucially, CAP uses a two‑step chain‑of‑thought prompt—first asking the model to describe the video, then to reason about any physics issues—so it does not assume videos are incorrect by default and yields only minimal false‑positive increases (0.001 in Semantic Adherence, 0.014 in Physical Commonsense), confirming its neutrality. Moreover, because it operates solely on the final video and its physics‑based standards, CAP is agnostic to the underlying generative pipeline—whether text‑to‑video, image‑to‑video, video‑to‑video, or hybrid.
@@ Appendix D Detailed Analysis on Evaluator
Appendix E Model Application Details
@@
Appendix F CAP Results
Here we further include the automatic evaluation results of CAP on the video generation models on Tab. 7.
Table 7: Automatic Evaluation results from CAP on proprietary and open-source video models.
Model SA PC Both
Proprietary Models
Pika 2.0 0.354 0.250 0.205
Luma 0.275 0.177 0.135
Kling-1.6 0.343 0.237 0.177
Sora-Turbo 0.357 0.225 0.176
Gen-3 0.380 0.164 0.133
Open-sourced Models
Hunyuan 720p 0.323 0.230 0.183
Step-Video-T2V 0.335 0.187 0.153
Open-Sora-Plan 1.3 0.143 0.110 0.065
Open-Sora 2.0 0.330 0.196 0.143
CogVideoX-1.5-5B 0.303 0.187 0.129
Wanx-2.1 0.365 0.243 0.191
LTX-Video 0.197 0.075 0.052
Appendix G Insights into Performance Between Models
Our analysis reveals that both model size and sampling budget significantly impact physics performance. Larger models consistently outperform smaller ones on foundational physics tasks, as shown in Table 8. For instance, the fact that Cosmos 14B outperforms Cosmos 7B across all SA/PC metrics confirming that scaling yields clear gains in semantic adherence and physical commonsense.
@@ Appendix G Insights into Performance Between Models
Cosmos 14B Agarwal et al. (2025) 0.374 0.268 0.212 0.331 0.243 0.182 0.091 0.069 0.020 0.333 0.241 0.184
Cosmos 7B Agarwal et al. (2025) 0.354 0.243 0.201 0.339 0.221 0.185 0.088 0.054 0.020 0.323 0.218 0.178
CogVideoX-1.5-5B Hong et al. (2022) 0.419 0.266 0.192 0.308 0.207 0.154 0.062 0.021 0.000 0.351 0.225 0.163
CogVideoX‑2B Hong et al. (2022) 0.382 0.247 0.175 0.246 0.199 0.148 0.073 0.021 0.010 0.310 0.210 0.150
@@
Appendix H Detailed Performance over All Physics Categories
In the main paper, we presented results across different levels of physics. Here, we provide a detailed breakdown of results for each specific physics category in Tab. 9.
@@ Appendix H Detailed Performance over All Physics Categories
CogVideoX-1.5 0.476 0.272 0.223 0.333 0.125 0.115 0.438 0.198 0.156 0.470 0.500 0.350 0.495 0.326 0.179 0.304 0.174 0.130 0.143 0.010 0.000 0.367 0.306 0.235 0.413 0.304 0.228 0.062 0.021 0.000
Step-video-T2V 0.401 0.265 0.204 0.273 0.202 0.166 0.361 0.239 0.145 0.371 0.264 0.227 0.384 0.273 0.238 0.328 0.227 0.154 0.165 0.000 0.000 0.388 0.385 0.257 0.446 0.368 0.316 0.085 0.021 0.020
@@ Appendix I Limitations
Appendix J Code and Data Availability
@@
Appendix K Query Used
Here we list the three MLLM queries used in this work, corresponding to the first-stage description prompt (Tab. 10), the second-stage structured-answer prompt (Tab. 11), and the detailed-prompt refinement query (Tab. 12). Note that for proprietary MLLMs such as GPT-o1, GPT-4o and Gemini-2.0-Flash, the model may refuse to evaluate a small fraction of generated videos; the refusal rate is below 3
@@
Appendix L Qualitative Analysis on Individual Model
@@
  • CogVideoX-1.5: Tends to produce distorted objects, such as non-spherical balls that appear dented. Movements are often excessively fast, blurred, erratic, or jittery, sometimes appearing unnaturally sped up.
@@
Figure 16: Physics Violations Observed in Videos Generated by the CogVideoX-1.5 5b Model. a. Raindrops do not return to the clouds after hitting the ground; instead, they exhibit normal rainfall, failing to demonstrate anti-physics. b. Toy car and real car rolling down ramps are depicted as the same size, violating proportional realism. c. Car driving on an icy road shows no slippage, unrealistically maintaining normal traction. d. Basketball dropped onto the ground displays unnatural bouncing motion. e. Person standing on one leg during yoga exhibits unrealistic human anatomy.
@@
Appendix M Static-Scene Rationalization as an Emerging Failure Mode
Beyond explicit physical violations, we observe that several text-to-video models exhibit a distinct failure mode in which the model generates a partially or fully static scene to avoid producing incorrect physical motion. Instead of attempting the intended physical interaction, the model suppresses motion entirely while still producing a visually coherent video. This behavior effectively “rationalizes” the violation by minimizing dynamics.
To systematically investigate this phenomenon, we analyzed 600 failure cases from our evaluation set, sampling 50 failed videos from each of the 12 models. Each video was annotated for whether it exhibits static-scene rationalization, defined as replacing the expected motion with a static or minimally animated scene.
Table 13: Rationalization rate among failure cases. Values represent the percentage of failed videos exhibiting static-scene rationalization.
Model Pika Sora Kling Luma Wanx Gen-3 Hunyuan Step CogVideo Open-S Open-S-P LTX
Rate 34% 24% 24% 22% 28% 14% 20% 18% 16% 4% 6% 4%
A clear trend emerges: models with higher overall generation quality tend to exhibit substantially higher rationalization rates. Lower-performing models generally fail due to more fundamental issues—such as incorrect object identities, distorted geometry, or inconsistent motion—leaving little opportunity to “rationalize.” In contrast, stronger models, once capable of producing semantically aligned and visually coherent scenes, often avoid explicit physics violations by freezing or minimizing motion, thus prioritizing aesthetic quality over physical correctness.
This pattern highlights a key challenge for future text-to-video systems. As models advance in visual fidelity, they may increasingly adopt conservative strategies that circumvent difficult physics rather than attempting them. Quantifying this emerging failure mode is essential for developing models that do not merely avoid errors but actively engage in physically grounded reasoning.
Appendix N Video Length from Models
Table 14: Video Length by Model
CogVideo Gen-3 Hunyuan Kling LTX-Video Luma Open-Sora-Plan Open-Sora Pika Sora Wanx Step-Video-T2V
Duration (s) 5 5 5 5 5 5 5 2 5 5 5 8
Appendix O Further Analysis on Generative Models Struggling with Certain Physical Scenarios
@@
Appendix P Leaderboard
Tab. 15 presents the leaderboard performance on both human evaluation and our designed CAP. The performance shows a highly consistency and only difference is that for proprietary models, Sora ranked number 2 while Kling ranked number 2 in CAP. Through further analysis, we found that Kling’s more cinematic style can bias CAP toward higher scores despite only moderate physical correctness, whereas our human annotators—who were instructed to focus strictly on physics plausibility—penalize these artifacts, causing Sora to rank higher under human evaluation.
@@
Appendix Q Video Quantitative Analysis
Here we present the quantitative analysis of the generated video.
Table 16: Quantitative analysis on Erratic Visual Changes.
Group Style Pass / Total Rate
All models Abrupt 53 / 240 22.1%
Non-abrupt 101 / 240 42.1%
Pika Abrupt 6 / 20 30.0%
Non-abrupt 15 / 20 75.0%
Table 17: Quantitative analysis on scene complexity.
Group Complexity Pass / Total Rate
All models Complex 28 / 240 11.7%
Less-complex 44 / 240 18.3%
Pika Complex 3 / 20 15.0%
Less-complex 8 / 20 40.0%
Appendix R Influence of Multiple Objects
@@
Appendix S Dataset Structure
@@
Appendix T Data Release
We publicly released a comprehensive dataset that includes the core assets of PhyWorldBench, the 1,050 example json file, the evaluation standards, and the physics categories and subcategories, used to assess physical realism in text-to-video models.